OpenStudy (anonymous):

Medal and Fan! A quadratic equation is shown below: 3x2 - 16x + 2 = 0 Part A: Describe the solution(s) to the equation by just determining the radicand. Show your work. (3 points) Part B: Solve 9x2 + 3x - 2 = 0 by using an appropriate method. Show the steps of your work, and explain why you chose the method used. (4 points) Part C: Solve 3x2 - 10x + 2 = 0 by using a method different from the one you used in Part B. Show the steps of your work. (3 points)

OpenStudy (anonymous):

@jim_thompson5910 could you help me please?

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

\[3x^2 - 16x + 2 = 0\] i assume by "radicand" they really mean "discriminant"

OpenStudy (anonymous):

\[b^2-4ac\] or in your case \[(-16)^2-4\times 3\times 2\]

OpenStudy (anonymous):

Okay so -232

OpenStudy (anonymous):

i get \(232\) a positive number that means there are "two real solutions"

OpenStudy (anonymous):

Right sorry I didnt take in account the negative times a negative

OpenStudy (anonymous):

So part a = 232 plus all the work shown and two real solutions :)

OpenStudy (anonymous):

9x^2+3x-2=0

OpenStudy (anonymous):

What is my first step to doing part b? :)

OpenStudy (anonymous):

me scrolling up to read the question!!

OpenStudy (anonymous):

\[9x^2 + 3x - 2 = 0 \] right?

OpenStudy (anonymous):

yes :D

OpenStudy (anonymous):

by some miracle this factors as \[(3x-1)(3x+2)=0\]

OpenStudy (anonymous):

set each factor equal to zero and solve, i.e. solve \[3x-1=0\] and then solve \[3x+2=0\]

OpenStudy (anonymous):

okies so 3x-1=0 add 1 to both sides 3x=1 divide 3 x=.33 repeating

OpenStudy (anonymous):

3x+2=0 subtract 2 by each side 3x=-2 divide by 3 x= .66

OpenStudy (anonymous):

lol put away your calculator dear

OpenStudy (anonymous):

2/3 we leave it in fraction form?

OpenStudy (anonymous):

okay x= 1/3 and x=-2/3

OpenStudy (anonymous):

right :3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay great! Part C: Solve 3x2 - 10x + 2 = 0 by using a method different from the one you used in Part B. Show the steps of your work. (3 points)

OpenStudy (anonymous):

What equation do we use this time :D

OpenStudy (anonymous):

use the quadratic formula \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with ' \[a=3,b=-10,c=2\]

OpenStudy (anonymous):

\[x=10+\sqrt{100-(-4)(3)(2)} \]

OpenStudy (anonymous):

which is all over 2(3) @satellite73 ?

OpenStudy (anonymous):

lets start in the square root, 100-(-4)(3)(2) = 100+24=124

OpenStudy (anonymous):

10+SQRT of 124 / 6= x

OpenStudy (anonymous):

that is a good start

OpenStudy (anonymous):

but the square root of 124 isnt a whole number

OpenStudy (anonymous):

true

OpenStudy (anonymous):

we got a mistake here i didn't catch

OpenStudy (anonymous):

\[\sqrt{b^2-4ac}=\sqrt{(-10)^2-4\times 3\times 2}=\sqrt{100-24}\]

OpenStudy (anonymous):

Oops! that makes it easier!

OpenStudy (anonymous):

so \[\frac{10\pm\sqrt{76}}{6}\]

OpenStudy (anonymous):

that is still not a whole number... @satellite73

OpenStudy (anonymous):

what do I do now is that the final answer?

OpenStudy (anonymous):

@satellite73 ?

OpenStudy (anonymous):

you can write it in simplest radical form if you like

OpenStudy (anonymous):

Thank you so much for all your help <3 :)

OpenStudy (anonymous):

\[\sqrt{76}=\sqrt{4\times 19}=2\sqrt{19}\]

OpenStudy (anonymous):

so\[\frac{10\pm2\sqrt{19}}{6}=\frac{5\pm\sqrt{19}}{3}\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

be careful when you cancel, you have to cancel from both terms

OpenStudy (anonymous):

and yes it is not a whole number, which is why you can't factor using integers